OpenStudy (anonymous):

Complete the equations of the system in slope-intercept form. Use a decimal for the slope if necessary. Line 1 x y –3 7 –2 5 Line 2 x y 0 5 2 1 Line 1: y = x + Line 2: y = x + @iGreen

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Okay, first we find the slope between the two points for Line 1. (-3, 7), (-2, 5) \(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(m = \dfrac{5-7}{-2+3}\) Solve that, tell me what you get.

OpenStudy (igreen):

Subtract 5 - 7 Add -2 + 3 What do you get? @igivemedalsandfollow!